OpenStudy (anonymous):

find lim h->0 [sin(6x+6h)-sin(6x)/h]

OpenStudy (anonymous):

hey ,is this the original q ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

is everything \[\frac{\sin(6x+6h)-\sin6x }{ h }\] like this?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Wow, that matches the definition of a derivative.

Parth (parthkohli):

Have you been taught derivatives or do you have to compute this limit without any help of those?

OpenStudy (anonymous):

THE DEFINATION OF THE DERIVATIVE???? DO U KNOW IT?

OpenStudy (anonymous):

yes but i was taught it a while ago and dont remember how to solve it. I know derivatives.

OpenStudy (raden):

yeah, the question like to find the derivative of y = sin(6x)

Parth (parthkohli):

OK, so this is the definition of the derivative of sin(6x). Do you know the Chain Rule?

OpenStudy (anonymous):

yes for sin(6x) it would be cos(6x) (6)

Parth (parthkohli):

Wonderful! 6cos(6x) should be enough.

hartnn (hartnn):

but maybe the question requires the use of limit definition to find the answer.

Parth (parthkohli):

Hmm, but if he was taught derivatives, then why take the longer route?

hartnn (hartnn):

why go to derivative definition when the question specifically ask for the limit. (the only reason can be that the question itself belongs to derivative topic)

hartnn (hartnn):

so ej, does this belong to derivative topic or limit ?

Parth (parthkohli):

OP knows what a derivative is. That should be enough for him.

hartnn (hartnn):

ok, if ej doesn't mind, then i'm ok too :)

Parth (parthkohli):

Err, sorry... I might have appeared downright rude there

OpenStudy (anonymous):

It is a multiple choice questions the answer options are a.) 0 b.) 6sin(6x) c.)6cos(6x) d.) -6cos(6x) e.)-6sin(6x)

OpenStudy (anonymous):

i need to show work though

Parth (parthkohli):

Then our work is correct.

OpenStudy (anonymous):

how do I get rid of h in the denominator

OpenStudy (anonymous):

Even though I know the easy way to solve this i have to show the long way with work to get full credit.

Parth (parthkohli):

Do you want to do it with the definition of a limit?

OpenStudy (anonymous):

solving using the definition of derivative is what it is asking

Parth (parthkohli):

Then we have done it. The definition of a derivative is lim h->0 [f(x+h)-f(x)]/h Here, f(x) is sin(6x)